# Data-Juicer Sandbox: A Comprehensive Suite for Multimodal Data-Model Co-development

### v1

@@ 3.1 Motivation and Design
To accommodate the multifaceted workflows intrinsic to data-model co-development, we design a versatile, user-friendly sandbox laboratory, as depicted in Figure 1. The laboratory incorporates a spectrum of components for activities such as data analysis, filtering, recipe optimization, model training, inference, and evaluation, all orchestratable via configuration files. The architecture is stratified into three tiers: bespoke end-to-end workflows, generic development behaviors, and foundational data-model development capabilities:
@@
  • Moreover, users can flexibly introduce or innovate components at the capabilities level (such as novel models, metrics, or data processing algorithms) and behavior hooks (like multidimensional data quality assessments and adaptive adjustments based on multiple probe outcomes) interchangeably.
@@
This design allows for streamlined configuration and reuse of established infrastructure. Importantly, it expedites the prototyping of data and model development solutions, integrating actionable and quantifiable capabilities for swift feedback derivation and informed decision-making. Detailed explorations of the lower tiers are provided in Section 3.2, while Section 3.3 showcases the sandbox’s utility in establishing a probe-refine-scaling workflow for data-model co-development.
3.2 Flexible Capability Factory and Behavior Hooks
Components for Advanced Data Processing.
Within the sandbox, we design numerous factory classes and behavior hooks for data processing and evaluation, simplifying and unifying the interfaces provided by the open-source Data-Juicer system (chen2024datajuicer). Users gain great flexibility to leverage over 100 feature-rich operators and dedicated tools for efficient data analysis, processing (including filtering, modifying, generating, etc.), and evaluation:
  • From an actionable perspective, key components include DataProcessors and DataRecipeRefiners. DataProcessors invoke off-the-shelf data processing operators from Data-Juicer, such as Filters and Mappers, enabling accelerated and scalable data processing through system optimization and parallel support. DataRecipeRefiners offer tools to refine recipe configurations, such as adjusting filtering thresholds based on k-sigma principle or percentile distribution to differentiate data subsets.
  • From a measurable perspective, DataAnalyzers and DataEvaluators are provided. DataAnalyzers encapsulate various data analysis tools from Data-Juicer, performing efficient computations of target metrics like text perplexity or video aesthetic value, and providing common statistical values (mean, variance, percentile thresholds). DataEvaluators include tools for evaluating data quality, such as GPT quality classifier.
Components for Streamlined Model Advancement.
@@ 3.2 Flexible Capability Factory and Behavior Hooks
Similarly, model development components integrate state-of-the-art open-source tools and libraries, streamlining interfaces and enhancing usability for rapid and user-friendly development experiences. Actionable components encompass training functionalities, such as Mini-Gemini (Li et al., 2023) for image-to-text, EasyAnimate (Xu et al., 2024a) for text-to-video, and ModelScope  [^1] https://github.com/modelscope/modelscope for general generative models. Measurable components offer evaluation interfaces like FVD (Unterthiner et al., 2019), VBench (Huang et al., 2024) for synthesized video assessment, GPT-4 (OpenAI, 2023) for generated text scoring and ranking, and TextVQA (Singh et al., 2019), MMBench (Liu et al., 2023), and MME (Fu et al., 2023) for image-to-text synthesis evaluation.
Model integration is facilitated through the use of subprocesses to call bash scripts of integrated libraries, with code and parameter adjustments ensuring a cost-controlled environment where, for example, a single GPU card can complete a training session within one day, providing rapid feedback.
Figure 2: A probe-analyze-refine workflow within the sandbox for systematic data-model co-development. (1) The process begins with cost-effective, contrastive explorations on small data pools to identify the most effective data processing operators (OPs). (2) These OPs are then combined and scaled within data recipes through hierarchical data pyramids to assess their cumulative impact on model performance. (3) Lastly, data utilization is optimized via duplication and diversity analysis, guiding the final formulation of refined data recipes and datasets for enhanced large-scale training. This structured workflow enables informed decision-making and efficient optimization for developing superior models.
3.3 A Probe-Analyze-Refine Workflow for Data-Model Co-Development
To demonstrate the usage of the sandbox laboratory, here we elucidate a built-in systematic workflow. As illustrated in Figure 2, we commence with a series of cost-effective contrastive explorations aimed at addressing several key questions, before judiciously applying them to enhance large-scale data-model co-development:
  1. 1.
    Initially, we seek to ascertain which data processing operators (OPs) contribute most effectively to enhancing specific target model metrics? This involves creating equal-size data pools, each processed uniquely by a singular OP and subsequently ranked by data metrics. Models are trained on these data pools, enabling us to perform an in-depth analysis of OP effectiveness and its correlation with model performance across various quantitative and qualitative indicators.
  2. 2.
    Guided by insights derived from the most impactful OPs that ranked highest by model metrics, we proceed to study whether these OPs can be effectively combined into data recipes and scaled up. To facilitate this exploration, we establish a hierarchical data pyramid, wherein data pools are categorized across different tiers based on their ranked model metric scores. This stratification also elucidates the viability of OP combinations when scaled with increased data volumes.
  3. 3.
    Finally, we delve into optimizing data utilization through a dual analysis focusing on duplication and diversity. We will assess whether the training process would benefit more from the repeated use of high-quality data pools or from the inclusion of lower-quality data to expand the overall data pool. Through a systematic examination of cost-controlled experiments, we can formulate optimized data recipes and datasets, which are then leveraged in more resource-intensive training sessions to cultivate models with superior performance.
3.3.1 Single-Operator Data Pools
Data-Juicer provides a rich assortment of OPs tailored for multimodal data encompassing text, images, videos, and audio. Given the variability in data sources, models, and downstream tasks, it is imperative to initially ascertain the impact of each OP to guide the optimization of data-model co-development effectively.
To this end, starting with an initial dataset $\mathcal{D}$, we define a single-operator data pool $\mathcal{P}_{i}$ as the dataset processed exclusively by the $i$-th OP ($OP_{i}$) available in Data-Juicer as follows:
$\mathcal{P}_{i}=\mathcal{DJ}[\mathcal{OP}_{i}(\rho_{i})](\mathcal{D}),$ (1)
where $\mathcal{DJ}$ denotes the data-processing function implemented by Data-Juicer, and $\rho_{i}$ is the hyper-parameters governing the operation of $\mathcal{OP}_{i}$. The OPs of Data-Juicer can compute specific statistics for every data pools, and apply threshold criteria to selectively filter or modify the data based on these statistics. Within this workflow, the initial dataset $\mathcal{D}$ into three equal-sized segments for filter OPs: $\mathcal{P}_{\text{high}}$, $\mathcal{P}_{\text{middle}}$ and $\mathcal{P}_{\text{low}}$, representing data with high, moderate, and low statistical measures, respectively. A randomly sampled $\mathcal{P}_{\text{random}}$ serves as a control group. This stratification fosters discriminative insights across varying degrees of data processing intensity.
Subsequently, models are trained independently on each data pool, undergoing comprehensive evaluation across multiple performance metrics. Throughout the training, we uphold consistent hyper-parameters and ensure $\mathcal{P}_{\text{random}}$ is of substantial size to yield a reliable and robust average performance across all downstream tasks. This setup enables the identification of best-performing OPs that universally excel or excel under specific evaluation criteria.
Importantly, we adhere to a cost-conscious design, implementing strategies to curtail expenses, such as downsizing the data pools, limiting training iterations, or adopting efficiency-enhancing techniques like LoRA (Hu et al., 2021). It enables the entire experimental pipeline to conclude at an affordable level. Specifically, all of our experiments can be completed within a single day utilizing a single A100 GPU, remaining both feasible and scalable for broader adoption in data-model co-development endeavors.
3.3.2 Multi-Operators Data Pools
Having explored the individual impact of each OP, our next step logically progresses to understanding the dynamics when multiple OPs are sequentially applied as in a data “recipe”. Our aim is to discern whether these OPs complement or counteract each other’s effects. To achieve this, we extend the data pool construction methodology from previous individual OP scenarios to a sequence of OPs:
$\begin{split}\mathcal{P}&=\mathcal{DJ}[\mathcal{OP}_{1}(\rho_{1}),\mathcal{OP}_{2}(\rho_{2}),...,\mathcal{OP}_{i}(\rho_{i})](\mathcal{D})\\ &=\mathcal{DJ}[\mathcal{OP}_{1}(\rho_{1})](\mathcal{D})\cap\mathcal{DJ}[\mathcal{OP}_{2}(\rho_{2})](\mathcal{D})\cap\cdot\cdot\cdot\cap\mathcal{DJ}[\mathcal{OP}_{i}(\rho_{i})](\mathcal{D}).\end{split}$ (2)
The number of possible multi-OP data pools grows exponentially with the addition of each new OP, necessitating a strategic selection of combinations to explore. Drawing on insights from single-OP experiments, we propose a pragmatic strategy: combining “Top” OPs, those with progressively diminishing impacts on model performance, while incrementally increasing the number of combinators.
Furthermore, we integrate the analysis of inter-OP relationships into our recipe formulation. Our workflow accommodates two approaches:
@@ 3.3.2 Multi-Operators Data Pools
  • The first method entails computing the Pearson correlation coefficients between the statistics generated by these OPs. Using a hierarchical clustering algorithm, we group the OPs into $k$ clusters. From each cluster, we select the OP with the strongest performance to form potential combinations.

### v2

@@ B.1 Overview Architecture
To support co-development based on feedback from data and model, we design the layered sandbox laboratory as illustrated in Figure 7. The laboratory incorporates a spectrum of components for activities such as data analysis, filtering, recipe optimization, model training, inference, and evaluation, all orchestratable via configuration files. The architecture is stratified into three tiers: bespoke end-to-end workflows, generic development behaviors, and foundational data-model development capabilities.
@@
  • Moreover, users can flexibly introduce or innovate classes at the capabilities level (such as novel models, metrics, or data processing algorithms) and behavior hooks (like multidimensional data quality assessments and adaptive adjustments based on multiple probe outcomes) interchangeably.
@@
This design allows for streamlined configuration and reuse of established infrastructure. Importantly, it expedites the prototyping of data and model development solutions, integrating actionable and measurable capabilities for swift feedback derivation and informed decision-making. Concrete explorations of the lower tiers are provided in Section B.2, while Section 3.2 showcases the sandbox’s utility in establishing a probe-analyze-refine workflow for data-model co-development.
B.2 Flexible Capability Factory and Behavior Hooks
Actionable Perspective.
Within the sandbox, we have created various factory classes and behavior hooks for data processing. These simplify and unify the interfaces offered by the open-source system, Data-Juicer. Users can utilize over 100 OPs and tools for data analysis, filtering, and synthesis. The toolkit automatically speeds up and scales up the processing through system optimization and parallel support. Classes for refining data recipes leverage tools like adjusting percentile distributions to distinguish data subsets or applying k-sigma rules to remove outliers.
@@ B.2 Flexible Capability Factory and Behavior Hooks
Besides, model development classes integrate SOTA open-source libraries, streamlining interfaces and enhancing usability for rapid and user-friendly development experiences. For example, models can be easily trained with diverse functionalities, such as Mini-Gemini for image-to-text, EasyAnimate and T2V-Turbo for text-to-video, and ModelScope for general generative models.
Measurable Perspective.
The sandbox also provides many classes to encapsulate observational capabilities, enabling subsequent optimization. For example, Data-Juicer-based classes can efficiently compute metrics such as text perplexity, video aesthetic value, and image quality scores with GPT API calls, along with statistical values like mean, variance, and percentiles. For models, various evaluation benchmarks are supported, like VBench and FVD (Unterthiner et al., 2019) for synthesized video assessment, and TextVQA (Singh et al., 2019), MMBench (Liu et al., 2023), and MME (Fu et al., 2023) for image-to-text evaluation.
More detailed developmental information can be found with the source code and accompanying documentation in our open-sourced link, which are also continuously maintained online and evolved. For example, users can conveniently run the following one-line entry script to easily run the sandbox and flexibly adjust configurations to switch between many built-in hooks and capabilities.
python tools/sandbox_starter.py --config configs/demo/sandbox.yaml
B.3 Extensibility of the Sandbox Suite
As a middleware, the sandbox itself does not impose any additional specific hardware dependencies. Instead, it inherits the dependencies of the integrated underlying libraries/frameworks and is continuously maintained to incorporate more community efforts in model training and evaluation. To simplify dependencies and avoid redundancy in an “all-in-one” environment, we have introduced and employed a lazy-loader mechanism at the Python package level. The code and parameters are provided and adjusted to ensure a stable and cost-controlled experiment environment, where, for example, a single GPU card can complete an end-to-end workflow within one day, enabling rapid feedback.
Besides, the capabilities of its underlying libraries are still evolving. For example, the utilized Data-Juicer’s operators are not limited to data filters investigated on the main page; they now encompass a wide range of functionalities, including 100+ Mappers, Filters, Deduplicators, and Selectors. This diversity allows for research on various types of data processing utilities within the Sandbox. Since Mappers can be employed to examine the effects of data augmentation and editing on downstream model task performance, which we reserved further exploration beyond filters as future work.
Appendix C Analysis on the Costs of the Proposed Sandbox Workflow
In this section, we give a more detailed discussion on the factors presented in the main page Sec. 3.2.4.
C.1 Impact of Sampling Ratio $r$ for Data Pools
In this section, we give some theoretical discussion to demonstrate the rationality of why scaling data pools can work in a cost-effective way.
Following the notations defined in Sec. 3.2.4, the probability that the error introduced by our experiments on the small-scale data pools is greater than $\epsilon$ can be denoted as:
$\mathbb{P}[\Delta_{pool}-\mathbb{E}[\Delta_{full}]\geq\epsilon].$
It is worth noting that our sandbox aims to obtain effective feedback with minimal expenditure. This feedback is often less apparent than what is obtained from the full dataset and larger model, meaning that $\mathbb{E}[|\Delta_{full}|]\geq\mathbb{E}[|\Delta_{pool}|]$. For operators that have a positive effect, we have $\mathbb{E}[\Delta_{full}]\geq\mathbb{E}[\Delta_{pool}]$. In this case, we can conclude that:
$\mathbb{P}[\Delta_{pool}-\mathbb{E}[\Delta_{full}]\geq\epsilon]$ $\leq\mathbb{P}[\Delta_{pool}-\mathbb{E}[\Delta_{pool}]\geq\epsilon]$
$\leq e^{-2\epsilon^{2}/(b-a)^{2}}.$
Here we use Hoeffding’s inequality with the assumption that $\Delta_{pool}\in[a,b]$. As the sample rate $r$ increases, the variance of the improvement across different training trials decreases, which is positively related to $(b-a)^{2}$, and thus the right term decreases. In conclusion, the probability of the error exceeding $\epsilon$ decreases exponentially as $r$ increases.
C.2 Impact of the Number $m$ of Planned Small-Pool Experiments
In the proposed sandbox workflow as illustrated in Sec. 3.2, we choose to split each data pool after applying our studied Data-Juicer OPs into three buckets according to the ranks of their generated statistics. The number The number of buckets acts as a multiplying factor to determine the total number of the planned small-pool experiments $m$, and finally the total cost of our experiments. It reflects the trade-off between total cost, data intervention intensity (via operator stats), and the informativeness of model feedback (the metric changes $\Delta$ on interested tasks compared to the models trained on random sampled data pools). More buckets lead to greater statistical differences between buckets with different ranks (especially the first and last ones), strengthening attribution to data processing effectiveness. However, more buckets also reduce per-bucket data, increasing the risk of inadequate data for models to exhibit reasonable $\Delta$.
As a result, we do not aim for models to be “training done” or “converged” in the sandbox workflow experiments. Instead, we want to observe enlightening changes—positive or negative—in models after targeted data intervention versus random data sampling. In our early experiments, we tested bucket counts of [2,3,4,5] to evaluate whether a model trained on randomly sampled data could reasonably decrease loss after one epoch and show statistically significant changes on downstream tasks. Our findings indicate that three buckets are empirically good for the studied scenarios. Once determined, all controlled experiments are aligned to one complete epoch and matched to the random pool data size.
Appendix D Implementation Details of Sandbox Experiments
D.1 Strategies to Combine OPs According to Correlations
In addition to assembling the OPs with the overall best performance, we also incorporate an analysis of inter-OP relationships into our recipe formulation process. Our workflow accommodates two strategies, with specific applications detailed in Section E.4:
@@ D.1 Strategies to Combine OPs According to Correlations
  • The first method involves computing Pearson correlation coefficients between the statistics generated by these OPs. Using a hierarchical clustering algorithm (Ward Jr, 1963), we group the OPs into $k$ clusters. From each cluster, we select the OP whose data pool yields the strongest model performance to form potential combinations.